OpenStudy (anonymous):

Find the inverse laplace of e^(2 - 2s)/(s-1).

OpenStudy (abb0t):

split the terms into two. that might help identify what you have.

OpenStudy (abb0t):

\[e^2(\frac{ e^{-2s} }{ s-1 })\]

OpenStudy (anonymous):

no but its e^(2 - 2s)

OpenStudy (anonymous):

so it would be e^(2)/e^(2s)

OpenStudy (abb0t):

\[e^2(e^{-2s}) = e^{2-2s} \]

OpenStudy (anonymous):

oh okay hmmm

OpenStudy (abb0t):

You're basically rearranging your given into some form that you can recognize and apply the laplace formulas to. Hope that helps.

OpenStudy (anonymous):

ha its more complicated than that but thanks anyways :)

OpenStudy (abb0t):

then there's partial fractions.

OpenStudy (anonymous):

i just need to find the inverse laplace of e^(2 - 2s)/(s - 1)

OpenStudy (anonymous):

no theres not haha

OpenStudy (abb0t):

im trying to remember. damit. lol

OpenStudy (anonymous):

cause i got the first part using the shift function u(t - a)f(t - a) and it worked perfectly

OpenStudy (anonymous):

oh i got it!!!

OpenStudy (anonymous):

aren't exponentials usually left behind in the time domain?

OpenStudy (anonymous):

not in real time

OpenStudy (anonymous):

what did it work out to?

OpenStudy (anonymous):

f(t) = e^(t - 1)u(t-1) - e^(2)*e^(t - 2)*u(t-2)

OpenStudy (anonymous):

e^(t-1)u(t-1) comes from the part i already solved - e^(-s)/(s-1)

OpenStudy (anonymous):

you gone dun it

OpenStudy (anonymous):

too bad i dont understand what it means haha